# ACDZero: Graph-Embedding-Based Tree Search for Mastering Automated Cyber Defense

### v1

@@ I Introduction
Automated cyber defense (ACD) systems are designed to monitor network environments and execute corrective actions—such as host isolation, service restoration, decoy deployment, and credential rotation—with minimal human input [vyas2023automated, tang2025human]. While reinforcement learning (RL) and deep reinforcement learning (DRL) have demonstrated potential in training defense policies within simulated environments [nguyen2021deep, tang2023edge], the approaches often face the challenge of balancing exploration and exploitation [DBLP:conf/icml/DamPDW25, 10.1007/978-981-96-6599-0_2, li2025inspo] in complex networks with large decision/state spaces, and thus require an expensive amount of samples to learn a reasonable defense policy. As modern cyber-adversaries become more sophisticated [chen2025neurosymbolic, chen2025perception, hong2025poster] and leverage multi-step strategies. Existing solutions relying on RL and DRL struggle to capture multi-step look-ahead defense planning while also meeting the required sample efficiency for agile cyber defense.
@@
To bridge this gap, we propose ACDZero, a framework for learning automated cyber defense policies through Monte Carlo Tree Search (MCTS) and graph-based latent-space planning. More precisely, we leverage MCTS, which has demonstrated strong performance in multi-step reasoning and complex planning problems [swiechowski2023monte] such as mastering chess and board games [muzero, DBLP:conf/ecai/CzechBK24, silver2017masteringchessshogiselfplay]-to explicitly model the exploration-exploitation tradeoff in ACD and use statistical sampling to guide decision making. By building a dynamic model of the ACD problem (in a latent graph-embedding space as introduced later), each search consists of a series of simulated ACD games of defense action self-play to traverse a tree from root state to leaf state of the ACD game. Each simulation proceeds by selecting in each node/state a defense action with respect to a dynamically-updated upper confidence tree (UCT) bound to balance exploration-exploitation. The final game result of each rollout is then used to weight the nodes in the ACD game tree and to update the UCT bounds for learning a search policy with optimal rewards.
@@
Unique challenges arise from applying MCTS to a complex ACD environment like the CAGE Challenge 4 (CC4) [kiely2025cage], involving multiple subnets, dynamic communications, random initialization, and a large set of servers/hosts and defense actions. We note that structural heterogeneity and unknown environment dynamics due to partial observability lead to significant challenges in representing the dynamic observations and network states to learn a dynamic model in MCTS. This dynamism and uncertainty render standard fixed-length vector representations brittle and prone to failure when deployed across varying network topologies [vargas2021review, ccavucsouglu2024novel, li2025dual, li2025sfmdiffusion, zeng2024local, li2024crowdsensing]. To this end, ACDZero makes novel use of a Graph Neural Network (GNN) as an invariant structural engine to encode network entities as typed nodes and edges [guan2024graph, yan2023swingnn]. This representation serves as the foundation for a learned latent dynamics model, which enables MCTS to perform ”virtual” look-ahead simulations without access to the ground-truth simulator [muzero].
@@ II Backgrounds
Automated Cyber Defense. ACD has transitioned from traditional rule-based heuristics  [vyas2023automated] toward RL frameworks capable of discovering optimal defensive policies through autonomous environment interaction [nguyen2021deep, hammar2020finding, kunz2022multiagent, jiang2022intelligent]. However, the efficacy of RL in network security is often constrained by state representation. Early approaches predominantly utilized fixed-vector encodings, where network features are mapped to static indices [ridley2018machine, kiely2309autonomous, miehling2015optimal, xu2023cnn]. Such representations impose an implicit dependency on node ordering; because they lack permutation invariance, even a minor re-indexing of hosts results in disparate feature vectors for functionally identical network states, severely inhibiting generalization.
@@
These architectural limitations are particularly pronounced in high-fidelity benchmarks such as the CC4  [kiely2025cage]. Unlike static environments, CC4 features a stochastically initialized topology where the number of hosts (5–15 per subnet) and active services (1–5 per host) vary across episodes. This structural fluidity prevents agents from memorizing specific configurations and demands strategies that are robust to topological shifts. To mitigate this, recent research has pivoted toward graph-based representations [king2025automated, collyer2022acd]. By encoding the network as an attributed graph—where entities are nodes and relations are edges—defenders can leverage GNNs to achieve the permutation invariance necessary for reasoning about structural patterns across heterogeneous network configurations [guan2024graph].
@@
Monte Carlo Tree Search. MCTS is applied widely to solve planning problems through sequential decision-making  [mcts_survey, tangmalinzero]. A typical MCTS involves four phases, i.e., Selection to choose actions from candidates via UCB-style strategies [ucb1], Expansion to sample new candidate actions for existing nodes, Simulation to obtain the corresponding payoffs, and Backup to update the cumulated returns along the search path. We denote the state by $s$ and action by $a$. For each node $(s,a)$ in the tree, there are statistics including the estimated value $\Phi(s,a)$, visiting count $N(s,a)$.
@@
MuZero [muzero] is a classic MCTS framework learning an internal dynamics model, allowing it to perform tree-based planning without access to the environment’s ground-truth rules or a simulator. This framework is composed of three learnable components parameterized by $\theta$: (i) a representation function $s_{0}=h_{\theta}(o_{1},\dots,o_{t})$ that transforms observations into a latent state, (ii) a dynamics function $(s_{k},r_{k})=g_{\theta}(s_{k-1},a_{k})$ that predicts the next latent state and immediate reward, and (iii) a prediction function $\left(\mathbf{p}_{k},v_{k}\right)=f_{\theta}(s_{k})$ which outputs the policy prior and state value.
@@
During the MCTS process, MuZero navigates the search tree entirely within this learned latent space. Starting from the root node, it applies a variant of the predictor Upper Confidence Bound (pUCT) [muzero] to select actions: $a=\arg\max_{a\in\mathcal{A}}Q(s,a)+P(s,a)\frac{\sqrt{\sum_{b}N(s,b)}}{N(s,a)+1}c_{1}$ where $P(s,a)$ is the prior probability from the policy head and $c_{1}$ is a constant. Statistics along the search path are updated during the Backup phase using a cumulative discounted payoff $G_{t,k}=\sum_{\tau=0}^{l-1-k}\gamma^{\tau}\hat{r}_{k+1+\tau}+\gamma^{l-k}v^{l}$. The mean action value $Q(s,a)$ is then updated as the average of these bootstrapped returns: $Q(s,a)=\frac{N(s,a)Q(s,a)+G_{t,k}}{N(s,a)+1}$.
@@ III-C ACDZero
Our proposed algorithm, ACDZero , adopts a MuZero-like framework [muzero] for decentralized cyber defense. The core objective is to employ MCTS as a policy improvement operator that enhances a GNN-based PPO [ppo] agent. MCTS performs lookahead search in learned latent space, generating an improved policy $\pi_{\text{mcts}}$ and value estimates that are distilled into decentralized actor-critic networks.
@@
  • Selection: Starting from root node $s_{t,0}$, the agent traverses the tree by selecting actions balancing exploitation and exploration. We employ the pUCT rule [muzero] with Dirichlet noise. At each node $s$, the agent selects action $a^{*}$ according to:

### v2

# ACDZero: MCTS Agent for Mastering Automated Cyber Defense

@@ I Introduction
Automated cyber defense (ACD) systems are designed to monitor network environments and execute corrective actions—such as host isolation, service restoration, decoy deployment, and credential rotation—with minimal human input [34, 32]. While reinforcement learning (RL) and deep reinforcement learning (DRL) have demonstrated potential in training defense policies within simulated environments [24, 31], the approaches often face the challenge of balancing exploration and exploitation [10, 4, 21] in complex networks with large decision/state spaces, and thus require an expensive amount of samples to learn a reasonable defense policy. As modern cyber-adversaries become more sophisticated [6, 7, 13] and leverage multi-step strategies. Existing solutions relying on RL and DRL struggle to capture multi-step look-ahead defense planning while also meeting the required sample efficiency for agile cyber defense.
@@
To bridge this gap, we propose ACDZero, a framework for learning automated cyber defense policies through Monte Carlo Tree Search (MCTS) and graph-based latent-space planning. More precisely, we leverage MCTS, which has demonstrated strong performance in multi-step reasoning and complex planning problems [29] such as mastering chess and board games [26, 9, 28]-to explicitly model the exploration-exploitation tradeoff in ACD and use statistical sampling to guide decision making. By building a dynamic model of the ACD problem (in a latent graph-embedding space as introduced later), each search consists of a series of simulated ACD games of defense action self-play to traverse a tree from root state to leaf state of the ACD game. Each simulation proceeds by selecting in each node/state a defense action with respect to a dynamically-updated upper confidence tree (UCT) bound to balance exploration-exploitation. The final game result of each rollout is then used to weight the nodes in the ACD game tree and to update the UCT bounds for learning a search policy with optimal rewards.
@@
Unique challenges arise from applying MCTS to a complex ACD environment like the CAGE Challenge 4 (CC4) [15], involving multiple subnets, dynamic communications, random initialization, and a large set of servers/hosts and defense actions. We note that structural heterogeneity and unknown environment dynamics due to partial observability lead to significant challenges in representing the dynamic observations and network states to learn a dynamic model in MCTS. This dynamism and uncertainty render standard fixed-length vector representations brittle and prone to failure when deployed across varying network topologies [33, 3, 20, 19, 37, 22]. To this end, ACDZero makes novel use of a Graph Neural Network (GNN) as an invariant structural engine to encode network entities as typed nodes and edges [11, 36]. This representation serves as the foundation for a learned latent dynamics model, which enables MCTS to perform ”virtual” look-ahead simulations without access to the ground-truth simulator [26].
@@ II Backgrounds
Automated Cyber Defense. ACD has transitioned from traditional rule-based heuristics  [34] toward RL frameworks capable of discovering optimal defensive policies through autonomous environment interaction [24, 12, 18, 14]. However, the efficacy of RL in network security is often constrained by state representation. Early approaches predominantly utilized fixed-vector encodings, where network features are mapped to static indices [25, 16, 23, 35]. Such representations impose an implicit dependency on node ordering; because they lack permutation invariance, even a minor re-indexing of hosts results in disparate feature vectors for functionally identical network states, severely inhibiting generalization.
@@
These architectural limitations are particularly pronounced in high-fidelity benchmarks such as the CC4  [15]. Unlike static environments, CC4 features a stochastically initialized topology where the number of hosts (5–15 per subnet) and active services (1–5 per host) vary across episodes. This structural fluidity prevents agents from memorizing specific configurations and demands strategies that are robust to topological shifts. To mitigate this, recent research has pivoted toward graph-based representations [17, 8, 5]. By encoding the network as an attributed graph—where entities are nodes and relations are edges—defenders can leverage GNNs to achieve the permutation invariance necessary for reasoning about structural patterns across heterogeneous network configurations [11].
@@
Monte Carlo Tree Search. MCTS is applied widely to solve planning problems through sequential decision-making  [2, 30]. A typical MCTS involves four phases, i.e., Selection to choose actions from candidates via UCB-style strategies [1], Expansion to sample new candidate actions for existing nodes, Simulation to obtain the corresponding payoffs, and Backup to update the cumulated returns along the search path. We denote the state by $s$ and action by $a$. For each node $(s,a)$ in the tree, there are statistics including the estimated value $\Phi(s,a)$, visiting count $N(s,a)$.
@@
MuZero [26] is a classic MCTS framework learning an internal dynamics model, allowing it to perform tree-based planning without access to the environment’s ground-truth rules or a simulator. This framework is composed of three learnable components parameterized by $\theta$: (i) a representation function $s_{0}=h_{\theta}(o_{1},\dots,o_{t})$ that transforms observations into a latent state, (ii) a dynamics function $(s_{k},r_{k})=g_{\theta}(s_{k-1},a_{k})$ that predicts the next latent state and immediate reward, and (iii) a prediction function $\left(\mathbf{p}_{k},v_{k}\right)=f_{\theta}(s_{k})$ which outputs the policy prior and state value.
@@
During the MCTS process, MuZero navigates the search tree entirely within this learned latent space. Starting from the root node, it applies a variant of the predictor Upper Confidence Bound (pUCT) [26] to select actions: $a=\arg\max_{a\in\mathcal{A}}Q(s,a)+P(s,a)\frac{\sqrt{\sum_{b}N(s,b)}}{N(s,a)+1}c_{1}$ where $P(s,a)$ is the prior probability from the policy head and $c_{1}$ is a constant. Statistics along the search path are updated during the Backup phase using a cumulative discounted payoff $G_{t,k}=\sum_{\tau=0}^{l-1-k}\gamma^{\tau}\hat{r}_{k+1+\tau}+\gamma^{l-k}v^{l}$. The mean action value $Q(s,a)$ is then updated as the average of these bootstrapped returns: $Q(s,a)=\frac{N(s,a)Q(s,a)+G_{t,k}}{N(s,a)+1}$.
@@ III-C ACDZero
Our proposed algorithm, ACDZero , adopts a MuZero-like framework [26] for decentralized cyber defense. The core objective is to employ MCTS as a policy improvement operator that enhances a GNN-based PPO [27] agent. MCTS performs lookahead search in learned latent space, generating an improved policy $\pi_{\text{mcts}}$ and value estimates that are distilled into decentralized actor-critic networks.
@@
  • Selection: Starting from root node $s_{t,0}$, the agent traverses the tree by selecting actions balancing exploitation and exploration. We employ the pUCT rule [26] with Dirichlet noise. At each node $s$, the agent selects action $a^{*}$ according to: